OpenStudy (anonymous):

Cal 3 Problem:

OpenStudy (anonymous):

\[\rho=2\csc \phi \rightarrow \rho=\frac{ 2 }{ \sin \phi }\rightarrow \rho \sin \phi=2 \rightarrow \rho^2 \sin^2 \phi=4 \rightarrow \rho^2 - \rho^2 \cos^2 \phi=4 \rightarrow \] \[\rightarrow x^2+y^2+z^2-z^2=4 \rightarrow x^2+y^2=4\]

OpenStudy (anonymous):

Can anyone explain how the x^2, y^2, and z^2 happened?

OpenStudy (anonymous):

ro= sqrt(x^2+y^2+z^2) phi is so called polar angle. Angle that position vector makes with z axis. So ro^2cos^2 phi =z^2

OpenStudy (anonymous):

It is just the transformation from spherical to cartesian coordinates

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

\[\Large \rho^2=x^2+y^2+z^2\]

OpenStudy (anonymous):

writting it now to make sense of it

OpenStudy (anonymous):

ok i understand but how did the \[-\rho^2\cos^2\phi\] come about?

OpenStudy (anonymous):